OpenStudy (anonymous):

solve by any method x^3-15x=0

OpenStudy (anonymous):

ok let's do it together first switch -15x to the right and see what happens

OpenStudy (anonymous):

x^3=15x

OpenStudy (anonymous):

now cancel the x's what do you get

OpenStudy (anonymous):

the third power equals 15

OpenStudy (anonymous):

no! x^3 means x^2*x so you have x^2*x=15x now you remove the x from both sides so what is left?

OpenStudy (anonymous):

x^2=15

OpenStudy (anonymous):

YES! so the answer to the equation is?

OpenStudy (anonymous):

x=the square root of 15

OpenStudy (anonymous):

Solve by any method x^2-x-7=0

OpenStudy (anonymous):

use the quadratic formula -b+/-sqrt b^2-4ac